OpenStudy (anonymous):

Suppose that in one week, your company takes in $2,220 in sales at one of its cell phone stores. The price breakdown per phone is as follows: Cell phone model A4: $35 Smartphone Z20: $50 Suppose that a total of 51 phones were sold. this is the supposed problem with the method of solving in a graph

OpenStudy (anonymous):

you know that the revenue is 2220, which is the sum of all phones sold so: 2220 = 35*x + 50*y, with x='cell phones sold' and y='smartphones sold' now; there are 51 phones sold in total, so x+y=51 this gives us the 2 equations relevant to this problem: 2220 = 35*x + 50*y x+y=51 this can be solved in many ways, propably the easiest: x+y=51 => x=51-y 2220=35*x+50*y 2220=35*(51-y)+50*y 2220=35*51-35*y+50*y 2220=1785+15*y 15*y=2220-1785 15*y=435 y=435/15 y=29 x+y=51 x+29=51 x=51-29 x=22

OpenStudy (anonymous):

ok i got that part but i cant get how to graph it what do i insert for the graph to work